OpenStudy (anonymous):

Solve the triangle. B = 73°, b = 15, c = 10 OPTIONS : C = 39.6°, A = 67.4°, a ≈ 14.5 Cannot be solved C = 44.8°, A = 62.4°, a ≈ 14.5 C = 39.6°, A = 67.4°, a ≈ 20.3 And State whether the given measurements determine zero, one, or two triangles. C = 30°, a = 28, c = 14 One Zero Two PLEASE HELP

OpenStudy (anonymous):

@chme anything ?

OpenStudy (campbell_st):

well you can use the rule of sine to find angle C \[\frac{Sin(B)}{b} = \frac{Sin(C)}{c}... or ... \frac{\sin(73)}{15}= \frac{\sin(C)}{10}\] solve for C then angle A is found by angle sum of a triangle... and side a is found by the rule of sines